### v2

# Fast and Accurate Algorithms to Calculate Expected Modularity in Probabilistic Networks

Abstract
Modularity maximization is a widely used community detection technique for deterministic networks. However, little research has been performed to develop efficient modularity calculation algorithms for probabilistic networks. Particularly, it is challenging to efficiently calculate expected modularity when all possible worlds are considered. To address this problem, we propose two algorithms, namely $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$, partitioning the possible worlds based on their modularities to significantly reduce the number of probability calculations. We evaluate the accuracy and time efficiency of our algorithms through comprehensive experiments.
Keywords Modularity calculation $\cdot$ probabilistic networks $\cdot$ algorithms
1 Introduction
Network data is often associated with uncertainty, which can arise because of randomness of systems under study, or inaccuracy of measurements. The first type of uncertainty can be found for example in computer networks, where network links can be unreliable. Inaccuracy of measurements is ubiquitous, for example, when estimating interaction probabilities in protein networks [1, 2, 3] or assessing the existence of social relations in social networks [4, 5]. To address these challenges, uncertain networks can be modeled as probabilistic graphs, where edges are associated with probabilities representing the likelihood of their existence [6, 7, 8].
This work is motivated by the importance of commu- nity detection, that is the identification of sets of nodes better connected to each other than to nodes outside the set, in the presence of uncertainty. As for deterministic networks, the ability to identify communities has numerous applications across various fields where a network is an appropriate data representation and uncertainty is common. For instance, in the field of biology, finding communities in probabilistic networks is crucial to predicting protein complex memberships [9]. In the context of intermittently connected networks, community detection can be used to determine delivery subgraphs [10].
More specifically, this paper focuses on the important problem of calculating expected modularity in probabilistic networks. Modularity is a measure of the ratio of edges falling within groups minus approximately the expected number in an equivalent network with edges placed at random [11]. Higher values of modularity indicate better community structures. Therefore, despite some limitations of modularity as an objective function [12, 13, 14] , modularity optimization [15] has emerged as one of the most popular approaches for finding communities and core-periphery structures in deterministic networks.
However, using modularity with probabilistic networks is challenging, and there are only a few studies on how to extend modularity to this context. One way of extending modularity to probabilistic networks is to calculate it over all possible worlds [16], corresponding to all the permutations of the edges in the network, and calculate its expected value. This approach is accurate but computationally expensive as it involves calculating modularity $2^{|E|}$ times, where $|E|$ is the number of edges in the network. In [3], the authors first sample $K$ graphs from all possible worlds, then calculate modularity for each sampled graph and take the average value. This approach can be very fast, but prioritizes execution time over accuracy: one can speed up the execution by choosing a lower $K$, but at the cost of obtaining a potentially inaccurate result (due to insufficient sampling) without a guaranteed approximation error. Other ways of extending modularity to probabilistic networks include regarding edge probabilities as edge weights or setting a threshold to convert probabilistic networks to deterministic networks by removing edges with probability lower than the threshold. However, these approaches have limitations. Regarding probabilities as weights generally leads to wrong results, as we show in our experiments. Setting a suitable threshold is also a complicated problem as it requires prior knowledge of the network structure and the chosen threshold greatly affects the resulting modularity [17].
In fact, in Section 4 we experimentally evaluate all the aforementioned approaches, and show that there is no known simple and general way of computing the expected value of modularity other than enumerating and computing modularity in all possible worlds, without a risk of obtaining inaccurate results.
In this paper, to avoid computing modularity in $2^{|E|}$ possible worlds, we propose an intuitive method for expected modularity computation, partitioning all possible worlds based on their modularities so that for each partition we can (1) easily compute the expected value of modularity for that partition, and (2) quickly compute or estimate the probability of that partition.
Please note that while this work is motivated by the importance of community detection, here we focus on the foundational problem of computing expected modularity given an input clustering (that is, modularity computation) and not on how to use modularity to identify good clusterings (that is, modularity optimization). In this paper a clustering is always given as an input of the algorithm, and expected modularity is computed for the input clustering.
Our contribution. We use an expression for modularity calculation based on communities instead of nodes. We exploit this expression to design efficient algorithms for calculating expected modularity in probabilistic networks. Our algorithm reduces time complexity from exponential to polynomial compared to the brute-force approach. Different from e.g. sampling, whose accuracy depends on the number of samples, our method always returns an accurate result, and its execution time does not depend on the edge probability distribution. Thanks to its ability to return an accurate result in polynomial time, our method also allows us to evaluate the traditional approaches used to analyse probabilistic networks but not providing guarantees on their accuracy, that is, sampling, thresholding, and weighting, when used to estimate expected modularity. We also show that our method is robust with regard to different network structures, and study the factors influencing its running time.
Roadmap. The rest of the paper is organized as follows: after reviewing the background in Section 2, we propose our algorithms in Section 3. In Section 4, we present a thorough simulation-based experimental evaluation on random networks and real-world networks with different properties. We summarize and discuss our results in Section 5.
2 Background
2.1 Probabilistic networks and relevant measures
2.1.1 Probabilistic networks
Consider a probabilistic graph $\mathcal{G}=(V,E,p)$, where $V$ corresponds to the set of nodes in $\mathcal{G}$, $E$ represents the set of edges, and $p:E\rightarrow(0,1]$ is a function that assigns probabilities to edges. We use $e_{ij}$ to indicate the edge between nodes $i$ and $j$, $p_{ij}$ as a shorthand for $p({e_{ij}})$, and we assume that $|E|=m$.
One can consider a probabilistic graph as a generative model for deterministic graphs. A deterministic graph $G$ is generated by $\mathcal{G}$ connecting two nodes $i,j$ via an edge with probability $p_{ij}$. We write $G\subseteq\mathcal{G}$ to denote that $G$ is generated by $\mathcal{G}$, and we consider the edge probabilities independent to each other [18, 19]. We assume possible worlds semantics [16], which corresponds to the set $\{G=(V,E_{G})\}_{E_{G}\subseteq E}$ of all possible deterministic graphs generated by $\mathcal{G}$ with their associated probabilities. The probability of observing such a set is:
$\begin{split}Pr(G)=\prod_{e\in E_{G}}p(e)\prod_{e\in E\backslash E_{G}}(1-p(e)).\end{split}$ (1)
There are $2^{m}$ distinct graphs that can be generated by $\mathcal{G}$.
2.1.2 Modularity
Broadly speaking, one solution for community detection in deterministic networks is to maximize the modularity score, which rewards partitions which capture many of the edges within communities and penalizes partitions with few or large communities. Giving a community labeling of vertices x, the modularity for an undirected network is given by:
$\begin{split}Q=\frac{1}{2M}\sum_{i}\sum_{j}(A_{ij}-\frac{k_{i}k_{j}}{2M})\delta(x_{i},x_{j}),\end{split}$ (2)
where $M$ is the number of edges, $A$ is the adjacency matrix, $k_{i}$ represents the degree of node $i$, $x_{i}$ is the label of node $i$, and $\delta(x_{i},x_{j})$ is the Kronecker delta function, which equals 1 when its arguments are the same and 0 otherwise.
Modularity ($Q$) can also be expressed in the following form based on [15]:
$\begin{split}Q=\sum_{c\in\mathcal{C}}\frac{|e_{c}|}{(|e_{c}|+|e_{c,\bar{c}}|+|e_{\bar{c}}|)}-(\frac{2|e_{c}|+|e_{c,\bar{c}}|}{2(|e_{c}|+|e_{c,\bar{c}}|+|e_{\bar{c}}|)})^{2}\ ,\end{split}$ (3)
where $\mathcal{C}$ is the set of communities, $V_{c}$ is the set of nodes in community $c\in\mathcal{C}$, and
$\begin{split}e_{c}&=\{e_{ij}\ |\ i,j\in V_{c}\},\\ e_{c,\bar{c}}&=\{e_{ij}\ |\ (i\in V_{c}\land j\notin V_{c})\lor(j\in V_{c}\land i\notin V_{c})\},\\ e_{\bar{c}}&=\{e_{ij}\ |\ i,j\notin V_{c}\}.\end{split}$ (4)
An example of $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ for a community $c_{i}$ is presented in Figure 1.
Figure 1: Edge partitions used in the community-based definition of modularity, for a community $c_{i}$.
2.1.3 Expected modularity
A deterministic graph $G_{w}$ generated by $\mathcal{G}$ can be considered as one possible world. From Eq. 1, we know the probability of each possible world in $\mathcal{G}$. Let Q be the discrete distribution of $Q$ on the $2^{m}$ possible worlds. The expected value of Q is:
$\begin{split}E(\textbf{Q})=\sum_{w=1}^{2^{m}}Q_{w}Pr(w),\end{split}$ (5)
where $Q_{w}$ is the modularity value of deterministic graph $G_{w}$ (that is, possible world $w$).
2.1.4 Entropy ratio
The entropy [20, 21] $H(\mathcal{G})$ of a probabilistic graph $\mathcal{G}$ is defined as the joint entropy of its edges $H(e)$ for all $e\in E$. Here we use different entropies to represent the different levels of uncertainty of a network. The larger entropy, the more uncertainty. Due to the edge independence, the formula of entropy is $H(\mathcal{G})=-\sum_{i<j}p_{ij}\log p_{ij}-\sum_{i<j}q_{ij}\log q_{ij}$, where $q_{ij}=1-p_{ij}$. To easily compare graphs with different sizes, we normalize such entropy by dividing by the number of edges. We call this entropy ratio, where entropy ratio =$\frac{|\mathcal{H}(\mathcal{G})|}{m}$, and the range is [0,1].
2.2 Existing computation methods
2.2.1 Brute-force method
One brute-force method to calculate expected modularity of probabilistic graph $\mathcal{G}$ is by dividing such probabilistic network into $2^{m}$ possible worlds [10], calculating the modularity on every possible world and their expected value according to Eq. 5. Therefore, the time complexity of this method is at least exponential on the size of the graph.
2.2.2 Sampling method
We use a sampling method called Monte Carlo (MC) estimator, which is introduced in [22]. Specifically, the MC estimator first generates $\theta$ samples from the probability distribution over the possible worlds, then calculates modularity for each sampled graph and computes the average modularity value:
$\begin{split}Q=\frac{1}{\theta}\sum_{t=1}^{\theta}Q_{t}.\end{split}$ (6)
Once the parameter $\theta$ has been specified, creating samples from the ensemble is straightforward. The detailed algorithm is depicted in Algorithm 1.
input : The number of graphs to sample $\theta$; nodes $V$; edges $E$; edges probabilities $p$
output : Graph set $Z$
1 Initialize graph set $Z$;
2 for $z\leftarrow$ 1 to $\theta$ do
3       Create an empty graph $g$ only with nodes $V$;
4       for each pair of nodes $i,j\in E$ do
5             Draw a uniformly random number $r_{ij}\in[0,1]$;
6             if $r_{ij}\leq p_{ij}$, then add the edge ($i$,$j$) to graph $g$;
7
8      add $g$ to $Z$;
9
10return $Z$;
Algorithm 1 Sampling method
2.2.3 Threshold method
Thresholding is a simple method currently used in community detection in probabilistic networks. The purpose of using threshold methods is to change probabilistic networks to deterministic networks. After setting a threshold, we remove edges whose probability is lower than the threshold and consider the others as deterministic.
3 Proposed algorithms
In this section we introduce two algorithms to compute expected modularity, exactly and approximately. The main notation introduced in this section is summarized in Table LABEL:tab:notations.
Table 1: Summary of notations
probabilistic graph $\mathcal{G}$
deterministic graph $G$
number of nodes in $\mathcal{G}$ $n$
number of edges in $\mathcal{G}$ $m$
set of nodes in $\mathcal{G}$ $V$
set of edges in $\mathcal{G}$ $E$
edge between node $i$ and node $j$ in $\mathcal{G}$ $e_{ij}$
edge between node $i$ and node $j$ in $G$ $l_{ij}$
set of possible worlds in $\mathcal{G}$ $W=\{w_{1},w_{2},\dots,w_{2^{m}}\}$
probability of possible world $w$ $Pr(w)$
set of communities in $\mathcal{G}$ $\mathcal{C}=\{c_{1},c_{2},\dots,c_{k}\}$
number of communities in $\mathcal{C}$ $k$
set of nodes in community $c$ $V_{c}$
set of edge probabilities in $\mathcal{G}$ $P=\{p_{1},p_{2},\dots,p_{2^{m}}\}$
set of partitions of possible words in $\mathcal{G}$ $D=\{d_{1},d_{2},\dots,d_{s}\}$
3.1 Possible world partitioning for expected modularity
The Possible World Partitioning for Expected Modularity ($\mathrm{PWP}^{\mathrm{EMOD}}$) algorithm is a variation of the brute-force method where instead of iterating through all possible worlds we group them into partitions, so that inside each partition all possible worlds have the same value of modularity. Using the alternative definition of modularity in Eq. 3, expected modularity can be rewritten as:
$\begin{split}E(\textbf{Q})=\sum_{w=1}^{2^{m}}\sum_{c\in\mathcal{C}}Q_{c}^{w}Pr(w)\ ,\end{split}$ (7)
where $w$ indicates a possible world, $Pr(w)$ is the probability of possible world $w$,
$Q_{c}^{w}=\frac{|e_{c}^{w}|}{(|e_{c}^{w}|+|e_{c,\bar{c}}^{w}|+|e_{\bar{c}}^{w}|)}-(\frac{2|e_{c}^{w}|+|e_{c,\bar{c}}^{w}|}{2(|e_{c}^{w}|+|e_{c,\bar{c}}^{w}|+|e_{\bar{c}}^{w}|)})^{2}\ ,$
and $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ refer to $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ in possible world $w$.
After rearranging the two sums:
$\begin{split}E(\textbf{Q})&=\sum_{c\in\mathcal{C}}\sum_{w=1}^{2^{m}}Q_{c}^{w}Pr(w)\ ,\end{split}$ (8)
we can partition the $2^{m}$ possible worlds so that in each partition the values of $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are constant. By notating $T_{x}=\max(|e_{c}^{w}|)$, $T_{y}=\max(|e_{c,\bar{c}}^{w}|)$, and $T_{z}=\max(|e_{\bar{c}}^{w}|)$, we can write:
$\begin{split}E(\textbf{Q})=\sum_{c\in\mathcal{C}}\sum_{x=0}^{T_{x}}\sum_{y=0}^{T_{y}}\sum_{z=0}^{T_{z}}Q_{c}^{xyz}Pr(xyz)\ ,\end{split}$ (9)
where $xyz$ indicates the partition (that is, the set) of possible worlds where $|e_{c}^{w}|=x$, $|e_{c,\bar{c}}^{w}|=y$, and $|e_{\bar{c}}^{w}|=z$. When $x=y=z=0$, the corresponding graph is an empty graph and $Q_{c}^{xyz}=0$. $Q_{c}^{xyz}$ is the constant value of $Q_{c}^{w}$ in all possible worlds $w\in xyz$, that is:
$\begin{split}Q_{c}^{xyz}=\frac{x}{(x+y+z)}-(\frac{2x+y}{2(x+y+z)})^{2}\ .\end{split}$ (10)
Notice that computing expected modularity using Eq. 9 we only need to iterate over $|\mathcal{C}|T_{x}T_{y}T_{z}$ terms instead of $2^{m}$, and for each term the value of $Q_{c}^{xyz}$ can be computed in constant time. The probability of partition $xyz$ can be expressed as a product of probabilities, because $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are always disjoint sets:
$\begin{split}Pr(xyz)=Pr(|e_{c}^{w}|=x)Pr(|e_{c,\bar{c}}^{w}|=y)Pr(|e_{\bar{c}}^{w}|=z)\ .\end{split}$ (11)
These probabilities can be computed using the definition of Poisson Binomial distribution, and the probabilities on the edges in $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are,
$\begin{split}Pr(|e_{c}^{w}|=x)=\sum_{A\in F_{x}}\prod_{e\in A}p(e)\prod_{e\in e_{c}\backslash A}(1-p(e)),\end{split}$ (12)
$\begin{split}Pr(|e_{c,\bar{c}}^{w}|=y)=\sum_{A\in F_{y}}\prod_{e\in A}p(e)\prod_{e\in e_{c,\bar{c}}\backslash A}(1-p(e)),\end{split}$ (13)
$\begin{split}Pr(|e_{\bar{c}}|=z)=\sum_{A\in F_{z}}\prod_{e\in A}p(e)\prod_{e\in e_{\bar{c}}\backslash A}(1-p(e)).\end{split}$ (14)
Then the joint probability $Pr(xyz)$ is,
$\begin{split}Pr(xyz)=\prod_{\varrho\in\{x,y,z\}}\sum_{A\in F_{\varrho}}\prod_{e\in A}p(e)\prod_{e\in\delta_{\varrho}\backslash A}(1-p(e)),\end{split}$ (15)
$\delta_{\varrho}=\left\{\begin{array}[]{rcl}e_{c}&\mbox{for}&\varrho=x\\ e_{c,\bar{c}}&\mbox{for}&\varrho=y\\ e_{\bar{c}}&\mbox{for}&\varrho=z\end{array}\right.$
where $F_{\varrho}$ is the set of all subsets of $\delta_{\varrho}$ of size $\varrho$.
As an example of how the partitions are defined, consider the probabilistic graph in Figure 2(a). Figure 2(b) shows a tree whose leaves enumerate all partitions $xyz$ defined by community $c$. The two branches from the top triangle represent respectively all the possible worlds where $e_{c}^{w}$ contains no edges ($x=0$) and all the possible worlds where $e_{c}^{w}$ contains one edge ($x=1$), in this case edge $l_{12}$. As we move down the tree, the circles in the second level represent all the possible worlds where $e_{c,\bar{c}}^{w}$ contains no edges ($y=0$), one edge ($y=1$, in particular $l_{23}$ or $l_{24}$), and two edges ($y=2$, in this case both $l_{23}$ and $l_{24}$). The same with the diamonds in the third level representing possible worlds with a specific number of edges in $e_{\bar{c}}^{w}$. The squares at the bottom of the figure represent all partitions. Figure 2(c) shows all possible worlds in partition 7, where $|e_{c}^{w}|=0,|e_{c,\bar{c}}^{w}|=1,|e_{\bar{c}}^{w}|=2$.
$\mathrm{PWP}^{\mathrm{EMOD}}$ is detailed in Algorithm 2. In lines 6, 8 and 10, we use Eq. 12, Eq. 13 and Eq. 14 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
Computational complexity: The worst-case time complexity of $\mathrm{PWP}^{\mathrm{EMOD}}$ is $O(k|E|2^{|E|})$. We can break the analysis down into two parts.
Lines 3 to 10 concern the exact calculation of probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(2^{|E|})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $|E|$, so the overall complexity is $O(|E|2^{|E|})$.
Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(|E|^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $|E|$.
Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(k(|E|2^{|E|}+|E|^{3}))=O(k|E|2^{|E|})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
(a) A probabilistic graph with a community $c$.
(b) A tree enumerating all partitions of possible worlds defined by community $c$.
(c) All possible worlds in partition 7.
Figure 2: An example demonstrating how, for a given community, all possible worlds of a probabilistic graph can be partitioned so that $|e_{c}^{w}|,|e_{c,\bar{c}}^{w}|$, and $|e_{\bar{c}}^{w}|$ are constant inside each partition.
3.2 Approximate partitioning for expected modularity
In Eq. 15, $F_{\varrho}$ will contain $\frac{T_{\varrho}!}{(T_{\varrho}-\varrho)!\varrho!}$ elements, the sum over which is infeasible to compute in practice. Approximate Possible World Partitioning for Expected Modularity ($\mathrm{APWP}^{\mathrm{EMOD}}$) is an approximation variation of $\mathrm{PWP}^{\mathrm{EMOD}}$ where we use Discrete Fourier Transform (DFT) to approximate the probabilities for disjoint sets $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$ and $e_{\bar{c}}^{w}$ [23, 24],
$\begin{split}Pr(|e_{c}^{w}|=x)&\approx\frac{1}{T_{x}+1}\sum_{l=0}^{T_{x}}C^{-lx}\prod_{\alpha=1}^{T_{x}}(1+(C^{l}-1)p_{\alpha}^{x}),\end{split}$ (16)
$\begin{split}Pr(|e_{c,\bar{c}}^{w}|=y)&\approx\frac{1}{T_{y}+1}\sum_{l=0}^{T_{y}}C^{-ly}\prod_{\alpha=1}^{T_{y}}(1+(C^{l}-1)p_{\alpha}^{y}),\end{split}$ (17)
$\begin{split}Pr(|e_{\bar{c}}|=z)&\approx\frac{1}{T_{z}+1}\sum_{l=0}^{T_{z}}C^{-lz}\prod_{\alpha=1}^{T_{z}}(1+(C^{l}-1)p_{\alpha}^{z}).\end{split}$ (18)
The probability of partition $xyz$ is then:
$\begin{split}Pr(xyz)&\approx\prod_{\varrho\in\{x,y,z\}}\frac{1}{T_{\varrho}+1}\sum_{l=0}^{T_{\varrho}}C^{-l\varrho}\prod_{\alpha=1}^{T_{\varrho}}(1+(C^{l}-1)p_{\alpha}^{\varrho}),\end{split}$ (19)
where $C=exp(\frac{2i\pi}{T_{\varrho}+1})$, $i=\sqrt{-1}$, and $p_{\alpha}^{x}$, $p_{\alpha}^{y}$, $p_{\alpha}^{z}$ are, respectively, the probabilities of edge $\alpha$ in $e_{c}^{w},e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$.
By substituting Eq. 19 into Eq. 9, we obtain,
$\begin{split}E(\textbf{Q})&=\sum_{c\in\mathcal{C}}\sum_{x=1}^{T_{x}}\sum_{y=1}^{T_{y}}\sum_{z=1}^{T_{z}}Q_{c}^{xyz}Pr(xyz)\\ &\approx\sum_{c\in\mathcal{C}}\sum_{x=1}^{T_{x}}\sum_{y=1}^{T_{y}}\sum_{z=1}^{T_{z}}[\frac{x}{(x+y+z)}-(\frac{2x+y}{2(x+y+z)})^{2}]\\ &\times\prod_{\varrho\in\{x,y,z\}}\frac{1}{T_{\varrho}+1}\sum_{l=0}^{T_{\varrho}}C^{-l\varrho}\prod_{\alpha=1}^{T_{\varrho}}(1+(C^{l}-1)p_{\alpha}^{\varrho}).\end{split}$ (20)
$\mathrm{APWP}^{\mathrm{EMOD}}$ is also detailed in Algorithm 2. In this case, differently from $\mathrm{PWP}^{\mathrm{EMOD}}$, in lines 6, 8 and 10 we use Eq. 16, Eq. 17 and Eq. 18 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
input : A set of communities; Edges: $E$; Edge probabilities: $p$
output : Approximated expected modularity: $Sum$
1 $Sum=0$;
2 for $c\leftarrow 1$ to $k$ do
3       Initialize $P_{c}$, $P_{c,\bar{c}}$ and $P_{\bar{c}}$ as empty arrays;
4       Partition edges into $e_{c}$, $e_{c,\bar{c}}$ and $e_{\bar{c}}$;
5       for $x\leftarrow 0$ to $T_{x}$ do
6             Store $Pr(|e_{c}^{w}|=x)$ in $P_{c}[x]$;
7
8      for $y\leftarrow 0$ to $T_{y}$ do
9             Store $Pr(|e_{c,\bar{c}}^{w}|=y)$ in $P_{c,\bar{c}}[y]$;
10
11      for $z\leftarrow 0$ to $T_{z}$ do
12             Store $Pr(|e_{\bar{c}}^{w}|=z)$ in $P_{\bar{c}}[z]$;
13      for $x\leftarrow 0$ to $T_{x}$ do
14             for $y\leftarrow 0$ to $T_{y}$ do
15                   for $z\leftarrow 0$ to $T_{z}$ do
16                         $Q_{c}^{xyz}\leftarrow\frac{x}{(x+y+z)}-(\frac{2x+y}{2(x+y+z)})^{2}$ (Eq. 10);
17                         $Sum=Sum+Q_{c}^{xyz}P_{c}[x]P_{c,\bar{c}}[y]P_{\bar{c}}[z]$
18
19
20
return Sum
Algorithm 2 $\mathrm{(A)PWP}^{\mathrm{EMOD}}$
Computational complexity: The worst-case time complexity of $\mathrm{APWP}^{\mathrm{EMOD}}$ is $O(k|E|^{3})$. We can break the analysis down into two main parts.
Lines 3 to 10 concern the approximation of probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(|E|^{2})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $|E|$, so the overall complexity is $O(|E|^{3})$.
Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(|E|^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $|E|$.
Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(k|E|^{3})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
4 Evaluation
This section has three main objectives. First, we show that our algorithms are as accurate as the brute-force method , but are several orders of magnitude faster. Notice that the the brute-force method iterates over all possible worlds and thus gives us the correct result. Then, we use our approximation algorithm to evaluate the main current alternatives to estimate expected modularity: weighting, thresholding, and sampling. Finally, we study the behaviour of our algorithm when we vary the number of communities, the distribution of community sizes, and the type of input network.
All the experiments except the comparison with the brute-force method have been performed on a macOS system, with max CPU frequency 2.4GHz. The comparison with the brute-force method has been performed on a Linux system, with max CPU frequency 5GHz. Both systems have the same 32GB memory capacity. All the code is in Python and available at https://github.com/uuinfolab/Expected Modularity Calculation in Probabilistic Graph.
Datasets. In our experiments we use both synthetic datasets, to control and examine the properties of the data that may affect running time and accuracy of the tested algorithms, and real-world datasets, to demonstrate the practical use of our methods.
To generate synthetic datasets, we use Stochastic Block Model, Forest Fire Network (FFN), Barabási-Abert (BA), Small world (SW), and Erdős-Rényi (ER). The parameters used to generate specific random graphs are specified later in this section for each experiment.
As real datasets, we use the Enron email network, a Facebook network, a protein-protein interaction (PPI) network, and a collaboration network, summarized in Table 2. The data sources used to generate these networks have been commonly used in the literature on probabilistic networks. In particular, two datasets (Enron, PPI) have been directly taken from the literature, while for those not provided by the authors we obtained probabilistic networks from the same sources following similar procedures and reproducing similar network statistics. For the two datasets we reproduced and PPI dataset, to keep the original topology structure, we included all the nodes in selected clusters detected by modularity optimization and all edges between those nodes. In our experiments, these clusters are also used as input to compute expected modularity.
Enron: this dataset consists of emails sent between employees of Enron between 1999 and 2001. Nodes represent employees and there is an edge between two nodes if at least one email has been exchanged between them. The original dataset with edge probabilities is provided by the authors in [7].
Protein-protein interaction (PPI): the dataset contains nodes that represent the proteins, edges that represent the interactions between two proteins, and associated probabilities. This dataset is extracted from a protein database that directly provides interaction probabilities, as described in [1].
Facebook: the dataset is obtained from a weighted Facebook-like Forum network, that originates from an online community for students at the University of California, Irvine. The data, downloaded from toreopsahl.com, includes the users who sent or received at least one message. Probabilities are computed using the method proposed in [25]. Specifically, we applied an exponential cumulative distribution function (CDF) of mean 2 to the weights. We sample nodes and edges from the original networks.
[…]
Name Nodes Edges Clusters
Enron 524 833 3
PPI 593 1185 5
Facebook 306 1217 4
Collaboration 523 1224 2
Table 2: Summary of real datasets.
4.1 Accuracy and execution time
To show the accuracy of our methods, we compare the values of expected modularity computed using $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$ with the true results, that we can obtain using a brute-force approach iterating through all the possible worlds. It is worth noting that $\mathrm{PWP}^{\mathrm{EMOD}}$ is not an approximation algorithm, and is expected to compute the exact same result as the brute-force method: the experimental results are provided to show the correctness of the mathematical derivations on which $\mathrm{PWP}^{\mathrm{EMOD}}$ is based. On the contrary, $\mathrm{APWP}^{\mathrm{EMOD}}$ computes an approximation of expected modularity with the aim of further reducing computing time. Therefore, this experiment also aims to inspect the accuracy of the approximation.
For the first experiment, we generate networks with three communities and varying sizes, using a stochastic block model (also known as planted-community structure network) [27], with the parameters shown in Table 3. To control the size of networks, we make $p_{in}$ and $p_{out}$ sometimes slightly different. For the second experiment, we generate random probabilistic graphs, using Erdős–Rényi model, with different numbers of edges and we fix the number of communities to 5. Table 4 shows that the results calculated by $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$ are always the same as the true values, calculated by the brute-force method, down to several significant digits (four in the table). We also observe that even for larger networks where we can no longer use the brute-force method, the results obtained by $\mathrm{APWP}^{\mathrm{EMOD}}$ is practically the same as the one produced by $\mathrm{PWP}^{\mathrm{EMOD}}$, which does not involve any approximations.
$m$ $n$ $k$ $\mathrm{nc}$ $p_{in}$ $p_{out}$
9 9 3 3 0.8 0.03
14 12 3 4 0.8 0.03
21 15 3 5 0.6 0.04
25 18 3 6 0.5 0.03
35 21 3 7 0.5 0.03
$m$ number of edges
$n$ number of nodes
$k$ number of communities
$\mathrm{nc}$ number of nodes in each community
$p_{in}$ edge density within communities
$p_{out}$ edge density between communities
Table 3: Experimental parameters.
$m$ $S^{\mathrm{PWP}}$ $S^{\mathrm{APWP}}$ $S^{BF}$
9 0.3900 0.3900 0.3900
14 0.4401 0.4401 0.4401
21 0.4257 0.4257 0.4257
25 0.4455 0.4455 0.4455
35 0.4720 0.4720
$m$ number of edges
$S^{\mathrm{PWP}}$ expected modularity score of $\mathrm{PWP}^{\mathrm{EMOD}}$
$S^{\mathrm{APWP}}$ expected modularity score of $\mathrm{APWP}^{\mathrm{EMOD}}$
$S^{BF}$ expected modularity score of brute-force method
Table 4: Expected modularity computed using the brute-force method, $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$.
Table 5 and Figure 3 show that both our methods are multiple orders of magnitude faster than the brute-force approach. Figure 3 adds larger networks to test the behavior of $\mathrm{APWP}^{\mathrm{EMOD}}$ where the two other methods cannot be used. Both $\mathrm{PWP}^{\mathrm{EMOD}}$ and the brute-force method have an exponential time complexity, although, $\mathrm{PWP}^{\mathrm{EMOD}}$ is significantly faster. Figure 3 confirms the polynomial time complexity of $\mathrm{APWP}^{\mathrm{EMOD}}$.
$m$ $T^{PWP}$ $T^{APWP}$ $T^{BF}$
9 0.0020 0.0001 0.1201
14 0.0082 0.0004 8.7528
21 0.1916 0.0004 2401.2359
25 3.0294 0.0006 59956.8956
35 226.5636 0.0030
$m$ number of edges
$T^{PWP}$ running time of $\mathrm{PWP}^{\mathrm{EMOD}}$
$T^{APWP}$ running time of $\mathrm{APWP}^{\mathrm{EMOD}}$
$T^{BF}$ running time of brute-force method
Table 5: Running time of brute-force , $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$.
Figure 3: Running time of brute-force, $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$, with log axes.
In the next sections we only use $\mathrm{APWP}^{\mathrm{EMOD}}$, given that it is both accurate and much faster than the brute-force method and $\mathrm{PWP}^{\mathrm{EMOD}}$.
4.2 Comparison with alternative methods
In this section we show that both weighting (that is, treating the probabilities as weights) and thresholding (that is, considering the network as a deterministic one after keeping only high-probability edges) lead to wrong estimations of expected modularity. On the contrary, sampling can be used to get more and more accurate results by increasing execution time, so we study how the balance between time and accuracy when using a sampling approach compares to our algorithm.
4.2.1 Weighting
In this experiment, we show that directly regarding probabilistic networks as weighted deterministic networks leads to a wrong expected modularity calculation. In particular, we first generate a base network using stochastic block modeling, where the parameters are $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$ From this base network, we generate multiple probabilistic networks by assigning probabilities to its edges. I n each probabilistic network we assign the same probability to all the edges: for example, in Figure 4, the network corresponding to the value 0.20 on the $x$ axis has all its edges having probability 0.20.
Figure 4 shows expected modularity computed using our algorithm (red) and the result obtained using weighted modularity. We can see that if we directly regard probabilistic networks as weighted deterministic networks, the final results are always the same for the different tested networks, and also different from expected modularity except for the case when the network is deterministic (where all edge probabilities equal 1).
Figure 4: Weighted and expected modularity calculated on the same graphs.
4.2.2 Thresholding
In this experiment, we compare expected modularity (as computed by $\mathrm{APWP}^{\mathrm{EMOD}}$) with the modularity obtained after thresholding, showing that the choice of threshold largely influences the final results, which are also not accurate in general.
Figure 5: Mean and expected modularity calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ and using the threshold method.
(a) CCS graph
(b) LCCS graph
(c) CCS graph
(d) LCCS graph
(e) CCS graph
(f) LCCS graph
Figure 6: Running time and expected modularity value calculated by different methods for graphs with a more clear (first row) and less clear (second row) modular structure, with different entropies.
Here we generate planted community structure networks with 110 edges and 3 communities. All networks have the same topology, generated by a stochastic block model with $p_{in}=0.72$ and $p_{out}=0.12$. We then generate different probabilistic networks by assigning random probabilities to the edges so that, for different networks, entropy ratio ranges from 0 to 1, and randomly assign those probabilities to the edges. For each network, we compute the modularities obtained using different thresholds (from 0.1 to 1). In Figure 5 we plot their mean (blue circles) and standard deviation (thick vertical lines). We can see that except for networks with low entropy ratios (that is, networks very close to deterministic networks), the modularities obtained through thresholding are far from the correct values (red line, calculated using $\mathrm{APWP}^{\mathrm{EMOD}}$).
4.2.3 Sampling
Thanks to the fact that using $\mathrm{APWP}^{\mathrm{EMOD}}$ we can compute expected modularity in polynomial time for a given input network, we can now test the ability of the sampling method to accurately estimate expected modularity, and study the factors influencing the balance between accuracy and execution time when sampling is used.
We use two kinds of random graphs, one is characterized by a pronounced community structure, we call it CCS graph, featuring dense intra-community edges and sparse inter-community edges; another exhibits a less distinct community structure, we call it LCCS graph. Both graphs contain 27 nodes and 3 communities, with similar number of edges by moving a few edges from within a community to between community from CCS graph to LCCS graph. The CCS graph and LCCS graph are constructed with stochastic block models where the parameters are, respectively, $k=3$, $\mathrm{nc}=9$, $p_{in}=0.99$, and $p_{out}=0.01$, and $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$. Notice that the CSS graph closely approximates three cliques. When the entropy ratio equals 1.00, all edge probabilities are assigned to 0.50; when the entropy ratio equals 0.47, all edge probabilities are assigned to 0.90; when the entropy ratio is 0.00, all edges are deterministic.
Figure 6 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. Specifically, the red nodes arranged in ’bands’ are the result of a different execution of the sampling method, all employing the same input $\theta$ as defined in Algorithm 1. For different executions, we have let the sampling method run for specific amounts of time (starting from the time needed by our algorithm), so that we could inspect the balance between execution time and likelihood to return an accurate result for the sampling method. For example, in Figure 6(f) we can see eight executions of the sampling method close to the blue vertical line: these were stopped after the same time used by our algorithm to return, and their $y$ value is the expected modularity they returned. From the figure we can see that, for different edge probability entropies, sampling always converges to the result produced by $\mathrm{APWP}^{\mathrm{EMOD}}$. Except for the case with no uncertainty (entropy ratio=0), sampling takes longer than $\mathrm{APWP}^{\mathrm{EMOD}}$ to converge toward the right results. By looking at Figures 6(b)6(d)6(f) or Figures 6(a)6(c)6(e), we find that when entropy ratio increases (larger uncertainty), sampling converges more slowly. If we compare two figures on the same column, e.g. 6(c) and 6(d), we find that when the graph structure is less clear, the convergence time of sampling is longer.
4.3 Factors influencing execution time
In Figure 7, we plot the execution time of $\mathrm{APWP}^{\mathrm{EMOD}}$ depending on the number of communities. For this experiment we generated a random probabilistic graph with 100 nodes and 500 edges. We vary the number of communities from 3 to 20, randomly assigning the nodes to them. From this figure, we observe that as the number of communities increases, the execution time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ also lengthens.
In Figure 8, we show the performance of $\mathrm{APWP}^{\mathrm{EMOD}}$ with the same number of communities but different community size distributions. From this figure, we can find that the larger variance among communities, the longer the running time.
Figure 7: Time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ with different numbers of communities.
Figure 8: Time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ with different variances of communities size.
In our final experiment, we examine the execution time of $\mathrm{APWP}^{\mathrm{EMOD}}$ on different random networks. In particular, we use 5 random graph models: a Forest Fire Network (FFN) [28], Barabási-Abert (BA) [29], Small world (SW) [30], Erdős-Rényi (ER) [31], and a clear community-structure graph (CCS graph). We use each of these models to generate three random graphs, all with around 200 nodes and 600 edges, but different numbers of communities. To provide a community assignment for the computation of modularity, we use a modularity optimization on the original deterministic graphs. We also control the number of communities, where $|C|=4$, $|C|=5$ and $|C|=6$ respectively.
Figure 9 shows different network structures containing 5 communities. Here nodes of the same color are in the same community computed using a modularity maximization method, and the number of colors represents the number of communities. Based on different network structures, we assign random edge probabilities so that entropy ratio is 0.4.
(a) Forest Fire network
(b) Barabási-Abert network
(c) Small world network
(d) Erdős-Rényi network
Figure 9: Networks generated by different models with 5 communities.
Figure 10 shows the running time of $\mathrm{APWP}^{\mathrm{EMOD}}$ on graphs with different network structures and numbers of communities. Generally, with an increasing number of communities the running time also increases, which fits the results in Figure 7. The visualization in Figure 9 suggests that a modularity optimization method may find a larger central community in FFN, which explains why the running time of FFN is always longer than for other networks. In fact, as shown in Figure 8, the larger the community size variance, the longer the running time.
Figure 10: Running time of our method for different models.
4.4 Real data
In this section, we apply $\mathrm{APWP}^{\mathrm{EMOD}}$ and the sampling method on real networks. Notice that we do not use thresholding and weighting, because the experiments on synthetic data have clearly shown that these approaches should not be used to compute expected modularity. Figure 12 shows different network structures with different numbers of clusters. Here nodes of the same color are in the same community computed using a modularity maximization method, and the number of colors represents the number of communities. Figure 11 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. For different executions, we have let the sampling method run for different amounts of time, so that we could inspect the balance between execution time and the likelihood of returning an accurate result for the sampling method. Figure 11 shows the same trends seen in the experiments with synthetic data. In these datasets, where the probability distributions are fixed and where clear community structures exist (we remind the reader that real data was obtained from high-modularity networks), the sampling method can produce accurate results in a short time.
(a) Enron network
(b) Facebook network
(c) PPI network
(d) Collaboration network
Figure 11: Running time and expected modularity value calculated by different methods in real datasets.
(e) Collaboration network $|E|=1224,|C|=2$
(f) PPI network $|E|=1185,|C|=5$
(g) Enron network $|E|=833,|C|=3$
(h) Facebook network $|E|=1217,|C|=4$
Figure 12: Real networks and the communities used in the experiments.
5 Discussion
In this paper, we present two methods for uncertain graph expected modularity calculation based on the partitioning of possible worlds.
When compared with a brute-force approach, both our methods are several orders of magnitude faster, with the approximate method reducing time complexity from exponential to polynomial without significantly affecting accuracy. For a network with 25 edges, which is close to the limit of what we could handle using exact methods before this paper, the execution time goes down from more than 15 hours (brute-force) to less than 5 seconds (our exact method) to a few microseconds (our DFT-based method). While this enables the usage of expected modularity in other algorithms, e.g. to identify communities, the execution time is however still high for large networks if the algorithm has to be repeated multiple times sequentially, taking several minutes for networks with thousands of edges.
Our method is in general preferable to other common approaches used to handle probabilistic networks. Weighting should not be used as it returns wrong results. The main reason is that node strengths (which equal node expected degrees in probabilistic networks) are multiplied in weighted modularity computation, while the expected product of degrees is not the same as the product of expected degrees in case of dependencies. Unfortunately, degrees in probabilistic networks are not independent, even when edge probabilities are, because edges are incident to two nodes and thus they either contribute to the degree of both or of none of them. A second reason is that weights are interpreted relative to each other in weighted modularity. So, for example, having all weights set to 0.01 or having them set to 1 makes no difference, while interpreting these numbers as probabilities means that in one case the network is probably empty, while in the second case all edges are almost certainly there.
In very specific cases, thresholding can be a fast alternative to our methods. However this is the case only when the network is almost deterministic. If we assume that all probabilities are at distance $\epsilon$ from 0 or 1, using thresholding we change edge probabilities from $1-\epsilon$ to 1 and from $\epsilon$ to 0, for all thresholds between $\epsilon$ and $1-\epsilon$. With small values of $\epsilon$, the change is small and the range of good thresholds is wide, making the method fast, accurate, and robust. However, as we are interested in handling probabilistic networks, it is not very useful to have a method that only works when the network is practically not probabilistic.
A similar consideration holds for sampling: sampling from a probabilistic graph with probabilities very close to 0 or 1 immediately gives an accurate result, the same computed on the thresholded deterministic network. We have also noticed how the likelihood of producing an accurate modularity using sampling increases when the community structure is very clear, close to a set of cliques. This is due to the fact that modularity depends on the number of edges inside communities, and not on the position of these edges. As a result, when many (and most) edges are inside communities, many sampled graphs will have a similar number of edges inside those communities, and thus similar values of modularity. Sampling from a narrow distribution has then a faster convergence. However, in most interesting cases our algorithm is preferable to sampling: it returns an accurate value whose accuracy does not depend on the execution time, while sampling can give an inaccurate result, in addition to not giving any indication of whether the result is accurate or not. The more the network is uncertain, the less well sampling works, while our algorithms are not affected by the probability distribution (neither with respect to accuracy nor execution time).
While the probability distribution does not impact the execution time of our method, the way in which nodes are assigned to communities does, both with respect to the number of communities and to the distribution of community sizes. The experiments on different types of networks (with or without a long-tail degree distribution, with or without a high clustering coefficient) also show that our algorithms are not significantly and directly affected by the network type, although depending on the network type it can be more or less likely to have to execute our algorithm with large communities, which would then indirectly affect execution time.
This paper opens the problem of how to use our algorithm as a sub-procedure of a community detection method. This problem adds two interesting aspects to our work: first, when used as a heuristic inside an optimization algorithm, a less accurate calculation of modularity may still lead to the same clustering, so there is a question about when it is needed to be accurate and when a faster approximation is sufficient. Second, we may have to execute the modularity calculation many times, which poses an additional computational challenge in particular for large networks. These two aspects may allow the usage of different computational approaches, for example to optimize a sequence of executions utilizing methods similar to the ones used in [32, 33].
Acknowledgments
This work has been partly funded by eSSENCE, an e-Science collaboration funded as a strategic research area of Sweden, by the Centre of Natural Hazards and Disaster Science, by the Wallenberg AI, Autonomous Systems and Software Program WASP, and by the AI4Research initiative at Uppsala University.